OpenStudy (anonymous):

can you help? Use guess, check and comment to find two numbers whose sum and product are given. Sum = 83, product = 1632

OpenStudy (angela210793):

form a system x+y=83--->x=83-y xy=1632 -----> (83-y)y=1632-->-y^2+83y-1632=0--->y=32 or y=51 so x=83-32=51 or x=83-51=32

OpenStudy (angela210793):

Get it?

OpenStudy (anonymous):

can you do it without forming a system? or a simpler way?

OpenStudy (anonymous):

guess...

OpenStudy (angela210793):

hmmmmmm...but this is a veryyyyyyyy easy way... 1st u form the system x+y=83 x*y=1632 in the first eq x=83-y and then we replace (83-y) instead of x then i guess u know how to solve this kind of eq

OpenStudy (anonymous):

true but not guess, check, comment....

OpenStudy (anonymous):

Guess 43 and 40. check product, comment yes/no too big/too small.. Next Guess, repeat, close in on answer.

OpenStudy (angela210793):

@Estudier O.o O.o :O

OpenStudy (anonymous):

Like developing an algorithm, low grade CS stuff:-)

OpenStudy (anonymous):

i will take your response into thought, thank you for answering, i am in the 9th grade and my teacher has obviously not taught me this yet :) what grade are you in?

OpenStudy (anonymous):

I'm in Europe and I am retired, I don't know what "9th grade" means..

OpenStudy (angela210793):

@Estudier ur in Europe too??? Where are u from?

OpenStudy (anonymous):

Born England, live in Spain.

OpenStudy (anonymous):

im australian, and 14

OpenStudy (anonymous):

So, like 3rd year of secondary school, something like that...

OpenStudy (angela210793):

Nice :)

OpenStudy (anonymous):

well thanks for the help. bye :)

OpenStudy (anonymous):

This question is in math class or computing?

OpenStudy (anonymous):

math class